# State Permutation Control in Non-Hermitian Multiqubit Systems with Suppressed Non-Adiabatic Transitions

### v1

@@ Abstract
Non-Hermitian systems have been at the focus of intense research for over a decade, partly due to their nontrivial energy topology formed by intersecting Riemann manifolds with branch points known as exceptional points (EPs). This spectral property can be exploited, e.g., to achieve controlled state permutations that are necessary for implementing a wide class of classical and quantum information protocols. However, the imaginary spectra of typical non-Hermitian systems lead to instabilities and breakdown of adiabaticity, which impedes the practical use of EP-induced energy topologies in quantum information protocols that explicitly rely on symmetric state flips. On the other hand, techniques that help to suppress non-adiabatic transitions in non-Hermitian systems have also been developed, but these advances have so far been limited to single qubits. In this work, we address this long-standing problem by introducing a model of interacting qubits governed by an effective non-Hermitian Hamiltonian that hosts EPs and possesses a completely real energy spectrum. We demonstrate that such non-Hermitian Hamiltonians enable realization of permutation groups in the multi-qubit eigenspace. Our findings indicate that, contrary to previous beliefs, non-Hermiticity can be utilized to achieve controlled state permutations in time-modulated multiqubit systems, thus paving the way for the advancement and development of novel quantum information protocols.
@@ I Introduction
However, subsequent studies revealed that symmetric state switching via dynamical encirclement of EPs is generally impractical due to the breakdown of adiabaticity in non-Hermitian systems, driven by the presence of an imaginary spectrum [17, 18, 19, 20]. In particular, any initial eigenmode of the NHH evolves into the state with the minimal loss through a series of non-adiabatic transitions (NATs) during the system dynamics. This intriguing chiral mode behavior has also been demonstrated in experiments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and various approaches for state control have been explored in the context of such quantum dynamics [33, *Ibanez_2014, 35, 36, 37, 38, 39, 40]. Although some progress has been made in restoring symmetric state flips to enable a programmable multi-state switch [41], the proposed approach does not fully recover the symmetry of the state transfer, as it still depends on the occurrence of NATs in the system.
@@
In a recent work in Ref. [42] some of us demonstrated that this seeming difficulty can be resolved in two-level systems by carefully choosing state trajectories in the system’s parameter space, allowing adiabaticity to be fully restored while dynamically winding around EPs. However, whereas the method can be potentially extended to linear optical multimode systems by combining the approach with findings in Ref. [41], it presents a significant challenge when applied to interacting multiqubit systems.
In this work, we address this long-standing problem by introducing an NHH that describes an interacting multiqubit system and exhibits a purely real spectrum with EPs defined on hyperboloid manifolds in the system parameter space. Remarkably, by dynamically encircling the EPs on such manifolds, it enables the implementation of a genuine and controllable symmetric multi-state switching; thus facilitating the generation of various permutation groups in the NHH eigenspace, a property that had been previously thought challenging or impossible to achieve.
@@
The ability to generate and control permutation groups in the NHH eigenspace holds promise for new opportunities in quantum information and communication. Potential applications include the production of diverse entangled states [43], the implementation of optimal quantum gates [44], the development of quantum error-correcting codes [45, 46], performing holonomic quantum computations [47], and advancements in quantum sensing [48], among others.
@@
As an illustrative example, we study here an NHH describing a three-qubit system and investigate the rich structure of eigenstate permutations that arise on dynamically orbiting its associated EPs. Our analysis reveals that the symmetric state switching protocol implemented here can generate a nontrivial non-Abelian permutation group comprising 576 elements. We also show that the proposed approach demonstrates robustness against external noise and perturbations, and it can be extended to any number of qubits. Our work therefore opens avenues for advancing and developing novel quantum information protocols, in which generating and steering permutations of quantum states plays a pivotal role.
@@ II Model
We start by introducing a tight-binding non-Hermitian Hamiltonian $\hat{H}\neq\hat{H}^{\dagger}$ describing $N$ qubits represented as spins with $ZZ$ coupling and which are subject to a transverse, in general non-uniform, effective complex magnetic field
$\hat{H}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}+\sum\limits_{m<n}J_{mn}\hat{\sigma}_{z}^{m}\hat{\sigma}_{z}^{n}.$ (1)
Here, the spin vector $\vec{{\boldsymbol{\sigma}}}=[\hat{\sigma}_{x},\hat{\sigma}_{y},0]^{T}$, with the symbol $T$ denoting the transposed operation, $\hat{\sigma}_{x,y,z}$ are the Pauli matrices, and the magnetic field vector is written as
$\vec{\mathbb{B}}_{||}^{k}=\Big{[}\alpha_{k}\sin\phi,\alpha_{k}\cos\phi,0\Big{]}^{T},$ (2)
@@
$\alpha_{k}=f_{k}\frac{y\sin\left({\rm Re}[\phi]\right)}{\sin\left({\rm Im}[\phi]\right)},\quad\phi=\arctan\left[(x+iy)^{-1}\right],$ (3)
with $f_{k},x,y\in{\mathbb{R}}$. For fixed $f_{k}$, the system parameter space is defined by a tuple $(x,y,J)$, with $J$ being the coupling matrix $J\equiv J_{mn}$.
The equations (2) and (3) determine the embedding map $(x,y)\to(B_{x}^{r},B_{x}^{i},B_{y}^{r},B_{y}^{i})$, where superscripts in $B$ denote real and imaginary parts of the magnetic field. The components of the local magnetic field, thus, describe a hyperboloid surface with a varied curvature $\alpha_{k}$ in ${\mathbb{R}}^{4}$ [42], namely
$(B_{x}^{r})^{2}+(B_{y}^{r})^{2}-(B_{x}^{i})^{2}-(B_{y}^{i})^{2}=\alpha_{k}^{2}.$ (4)
This parametrization of the magnetic field, together with the $ZZ$ qubit interactions, ensures that the NHH in Eq. (1) is pseudo-Hermitian and this symmetry is exact in the whole parameter space, meaning the spectrum of $\hat{H}$ is everywhere real. Indeed, the condition for the pseudo-Hermiticity of the NHH reads as [49]
@@
where $\eta$ is some Hermitian operator. Here, the operator $\eta$ can be chosen as a metric operator of the NHH describing non-interacting qubits, i.e., when $J=\boldsymbol{0}$ (see also Appendix A).
The NHH in Eq. (1) can be implemented in various experimental platforms, ranging from solid-state systems to photonic setups, utilizing either the Naimark dilation method [50] or photon postselection techniques [51]. We elaborate on this in more detail in Sec. V.
@@ III Eigenspectrum of the non-Hermitian Hamiltonian and ‘twisted’ exceptional points
Here we discuss the spectral properties of the NHH in Eq. (1), and the potential of the system with an odd number of qubits to realize a genuine permutation subgroup.
@@ III.1 Non-interacting qubits
We begin our analysis of the spectral properties of the NHH in Eq. (1) with the simplest case, when $J\equiv\boldsymbol{0}$, and $f_{k}=1$, for $\forall k$. In that case, the spectrum of $N$ non-interacting qubits is analogous to that of an $N$-dimensional hypercube [52, 53], namely
@@
which is generated by the spectrum of a single qubit $E^{\rm q}_{1,2}=\pm\alpha$. The latter we plot in Fig. 1. In what follows, we assume that the energy levels are always arranged in ascending order, i.e., $E_{1}\leq E_{2}\leq\dots\leq E_{2N}$.
Figure 1: Energy spectrum of a single qubit, described by the NHH in Eq. (1), as a function of the applied magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eqs. (6) and (3). The purely real-valued spectrum consists of two intertwined Riemann sheets with an EP located at $(x=0,y=1)$. Moreover, at this EP the spectrum of the NHH diverges, according to Eq. (8) (see the main text for details).
@@
where $\hat{R}=\exp\left(-i\pi\sigma_{x}/4\right)$ is a rotation operator. Because the pseudo-Hermiticity of the parameterized NHH does not experience spontaneous symmetry breaking, the left eigenvectors are simply the complex conjugates of the right eigenvectors [42].
We now proceed with a more detailed examination of the spectral properties of a single qubit, as its properties are directly reflected in Eq. (6). As can be seen from Fig. 1, the qubit spectrum is characterized by the presence of a singular point at $(x=0,y=1)$. Indeed, according to Eq. (3) one has
@@
where the sign at infinity depends on the direction of the limit approach to the singularity (see also Fig. 1). Furthermore, one finds the limit for the inner product of the normalized eigenvectors at this point to be
@@
Equation (9) implies that the qubit eigenvectors coalesce at this singular point. In other words, the singularity behaves like an EP, though it is not an EP in the usual sense that is encountered in systems with finite spectra. This seeming difficulty can be handled using mathematical tools borrowed from the projective space framework. Put differently, this singularity can be cast into an ordinary EP in some extended NHH, where the eigenvalues converge at the EP (see Appendix C for details).
@@
Nevertheless, for our purposes here, we can safely ignore the infinite nature of the spectrum at such EPs, because we are solely interested in the encircling trajectories around the EPs, along which both the energy spectrum and the system controlling parameters (magnetic fields and interactions) remain finite. These EPs may therefore be termed as ’twisted EPs,’ due to the nontrivial wrapping of the spectrum at these points that leads to the divergent behavior of the NHH eigenvalues. For simplicity, however, we will continue to refer to such singularities as EPs.
@@
When dynamically encircling the EP in Fig. 1 in the parameter space $(x,y)$, the eigenstates of a qubit can be interchanged adiabatically without invoking any non-adiabatic transitions [42]. This stems from the fact that the eigenvalues are everywhere real. Note that while a winding trajectory around the EP in the space $(x,y)$ can have the form of simple loops in a plane, the actual orbits in the vector space (${\mathbb{C}}^{2}$ or ${\mathbb{R}}^{4}$) of the magnetic field $\vec{\mathbb{B}}_{||}$ are the loops on a curved hyperboloid.
@@
We note that the spectrum described above drastically differs from that studied in Ref. [42], where the qubit eigenvalues have a well-defined limit in Eq. (8). The key distinction lies in the fact that, unlike in Ref. [42], the Riemann energy sheets are well separated at the point $(x=0,y=0)$ where the NHH becomes Hermitian (see Fig. 1). This is a very useful property, as it can allow performing state transfer protocols over multiqubit states which are often generated and exploited in closed systems. The latter thus enables the combination and extension of existing quantum information protocols, originally devised for Hermitian systems, to non-Hermitian systems as well.
@@
Figure 2: Eigenspectrum of the NHH $\hat{H}$ in Eq. (1) as a function of the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ for (a) two interacting qubits and (b) three interacting qubits. The system parameters are (a) $f_{2}=2f_{1}=2$, $J_{12}=1$; and (b) $f_{1,2,3}=1$, $J_{12}=0.5$, $J_{13}=1$; $J_{23}=1.5$. The spectrum consists of a stack of Riemann sheet pairs, whose number grows with the number of interacting qubits in the system. The set of EPs, constituting the exceptional line (EL), is located at the same point $(x=0,y=1)$ as in the case of a single non-interacting qubit in Fig. 1.

### v2

# Topological state permutations in time-modulated non-Hermitian multiqubit systems with suppressed non-adiabatic transitions

@@ Abstract
Non-Hermitian systems have been at the center of intense research for over a decade, partly due to their nontrivial energy topology formed by intersecting Riemann manifolds with branch points known as exceptional points (EPs). This spectral property can be exploited, e.g., to achieve topologically controlled state permutations that are necessary for implementing a wide class of classical and quantum information protocols. However, the complex-valued spectra of typical non-Hermitian systems lead to instabilities, losses, and breakdown of adiabaticity, which impedes the practical use of EP-induced energy topologies in quantum information protocols based on state permutation symmetries. Indeed, in a given non-Hermitian multiqubit system, the dynamical winding around EPs always results in a predetermined set of attenuated final eigenstates, due to the interplay of decoherence and non-adiabatic transitions, irrespective of the initial conditions. In this work, we address this long-standing problem by introducing a model of interacting qubits governed by an effective non-Hermitian Hamiltonian that hosts novel types of EPs while maintaining a completely real energy spectrum, ensuring the absence of losses in the system’s dynamics. We demonstrate that such non-Hermitian Hamiltonians enable the realization of genuine, in general, non-Abelian permutation groups in the multiqubit system’s eigenspace while dynamically encircling these EPs. Our findings indicate that, contrary to previous beliefs, non-Hermiticity can be utilized to achieve controlled topological state permutations in time-modulated multiqubit systems, thus paving the way for the advancement and development of novel quantum information protocols in real-world non-Hermitian quantum systems.
@@ I Introduction
However, subsequent studies revealed that symmetric state switching via dynamical encirclement of EPs is generally impractical due to the breakdown of adiabaticity in non-Hermitian systems, driven by the presence of a complex valued energy spectrum [17, 18, 19, 20]. In particular, any initial eigenmode of the NHH evolves into the state with the minimal loss (smallest imaginary energy component) through a series of non-adiabatic transitions (NATs) during the system dynamics, imbuing the dynamics with a chiral nature. This intriguing chiral mode behavior has also been demonstrated in experiments [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], and various approaches for state control have been explored in the context of such quantum dynamics [33, 34, 35, 36, 37, 38, 39, 40, 41].
The presence of NATs generically inhibits the possibility of using adiabatic dynamics to realize state–switching dynamics based on following eigensheets in the NHH spectrum. One possible approach to restoring state transfer is to employ shortcuts to adiabaticity (STA) [42, 41], where additional control fields are introduced to replicate the desired outcome of ideal adiabatic evolution. Alternatively, some progress has been made in restoring symmetric state flips to enable a programmable multi-state switch [43]. Both of these approaches presume that NATs appear in the system dynamics and are undesirable, and that one therefore seeks physical mechanisms to mitigate their effects.
@@
Here, we consider instead the approach of engineering systems in which non-trivial topological features appear in the NHH spectrum, but where NATs never disrupt the eigenstate evolution, thus avoiding the need for additional controls. That is, we consider here the realization of genuinely adiabatic multi-state transfer in a non-Hermitian context. In a recent work [44] some of us have demonstrated that such an approach is possible in two-level systems by carefully choosing state trajectories in the system’s parameter space on a submanifold described by a pseudo-Hermitian Hamiltonian with real eigenvalues, allowing adiabaticity to be fully restored while dynamically winding around EPs. However, while that method can potentially be extended to linear optical multimode systems by combining it with the findings in Ref. [43], it presents a significant challenge when applied to interacting multiqubit systems. Moreover, the method in Ref. [44] enables topological and adiabatic state transfer only among non-orthogonal eigenstates, since the entire system must be initialized in the non-Hermitian phase. This severely limits its applicability and integration with state-of-the-art quantum information protocols.
In this work, we address this long-standing challenge of adiabatic multi-state transfer in time-modulated non-Hermitian multiple qubit systems by introducing an NHH that describes an interacting multiqubit system and exhibits a purely real spectrum with specific EPs defined on hyperboloid manifolds in the system parameter space. At this special type of EPs, which we refer to as ‘twisted’ EPs, the Hamiltonian eigenvectors coalesce, although the corresponding eigenvalues diverge, with the sign of the divergence depending on the path taken to approach these EPs. Remarkably, by dynamically winding around such EPs along the Riemann eigensheets, we show that one can implement a genuine and controllable symmetric multi-state switching in the system dynamics, characterized by the absence of the NATs. This facilitates the generation of various permutation groups in the NHH eigenspace, a property that had been previously thought challenging or impossible to achieve due to the breakdown of the adiabatic theorem [45, 18, 19]. Moreover, the proposed NHH can also attain a Hermitian phase, where its eigenstates are orthogonal, and over which the multi-state switching protocol can also be performed. This allows to merge and combine existing quantum information protocols that have been developed for Hermitian systems, to non-Hermitian multi-qubit systems as well.
@@
The ability to generate and control permutation groups in the NHH eigenspace holds promise for new opportunities in quantum information and communication. Potential applications include the production of diverse entangled states [46], the implementation of optimal quantum gates [47], the development of quantum error-correcting codes [48, 49], performing holonomic quantum computations [50, 51, 52], and advancements in quantum sensing [53], among others.
@@
As an illustrative example, we study here an NHH describing a three-qubit system and investigate the rich structure of topological eigenstate permutations that arise while dynamically orbiting its associated EPs. These eigenstate permutations are topological because they do not depend on the shape of the winding loops around the EPs, and are determined solely by the intrinsic Riemann topology in the vicinity of the EPs. Our analysis reveals that the symmetric state switching protocol implemented in this work can generate a nontrivial non-Abelian permutation group comprising 576 elements. We also discuss the robustness of the protocol and its potential realization in various existing experimental setups. Our work therefore opens avenues for advancing and developing novel quantum information protocols using non-Hermitian quantum systems, in which generating and steering permutations of quantum states plays a pivotal role.
@@ II Model
We start by introducing a tight-binding non-Hermitian Hamiltonian $\hat{H}\neq\hat{H}^{\dagger}$ describing $N$ qubits represented as spins with $ZZ$ coupling, which are subject to a transverse, in general non-uniform, effective complex-valued magnetic field $\vec{\mathbb{B}}_{||}^{k}$:
$\hat{H}=\sum\limits_{k}^{N}\vec{{\boldsymbol{\sigma}}}^{k}\vec{\mathbb{B}}_{||}^{k}+\sum\limits_{k<l}J_{kl}\hat{\sigma}_{z}^{k}\hat{\sigma}_{z}^{l}.$ (1)
Here, the spin vector $\vec{{\boldsymbol{\sigma}}}=[\hat{\sigma}_{x},\hat{\sigma}_{y},0]^{T}$, with the symbol $T$ denoting the transpose operation, $\hat{\sigma}_{x,y,z}$ are the Pauli matrices, and the transverse magnetic field vector has $x,y$ components and is written as
$\vec{\mathbb{B}}_{||}^{k}=f_{k}\Big{[}\alpha\sin\phi,\alpha\cos\phi,0\Big{]}^{T},$ (2)
@@
$\alpha=\frac{y\sin\left({\rm Re}[\phi]\right)}{\sin\left({\rm Im}[\phi]\right)},\quad\phi=\arctan\left[(x+iy)^{-1}\right],$ (3)
with $f_{k},x,y\in{\mathbb{R}}$. For fixed $f_{k}$, the system parameter space is defined by a tuple $(x,y,\textbf{J})$, with J being the coupling matrix $\textbf{J}\equiv J_{kl}$.
The equations (2) and (3) determine the embedding map $(x,y)\to(B_{\rm x}^{r},B_{\rm x}^{i},B_{\rm y}^{r},B_{\rm y}^{i})$, where the superscripts in the four $B$ components denote the real and imaginary parts of the magnetic field in the $(\rm x)$ and $(\rm y)$ directions. The components of the local magnetic field thus describe a hyperboloid surface with a varied curvature $\alpha$ in ${\mathbb{R}}^{4}$ [44], namely
$(B_{\rm x}^{r})^{2}+(B_{\rm y}^{r})^{2}-(B_{\rm x}^{i})^{2}-(B_{\rm y}^{i})^{2}\equiv\alpha^{2}.$ (4)
This parametrization of the magnetic field, together with the $ZZ$ qubit interactions, ensures that the NHH in Eq. (1) is pseudo-Hermitian and that this symmetry is exact in the entire parameter space, meaning the spectrum of $\hat{H}$ is everywhere real. Indeed, the condition for the pseudo-Hermiticity of the NHH reads as [54]
@@
where $\eta$ is some Hermitian operator. Here, the operator $\eta$ can be chosen as a metric operator of the NHH describing non-interacting qubits, i.e., when $\textbf{J}=\boldsymbol{0}$ (see Appendix A).
The NHH of Eq. (1) with its complex-valued magnetic fields can be implemented in various experimental platforms, ranging from solid-state systems to photonic setups, by making use of either the Naimark dilation method [55] or photon postselection techniques [56]. We elaborate on possible experimental realizations in more detail in Sec. V.
@@ III Eigenspectrum of the non-Hermitian Hamiltonian and ‘twisted’ exceptional points
We now discuss the spectral properties of the NHH in Eq. (1), and the potential offered by this system to realize a genuine adiabatic permutation subgroup in non-Hermitian multiqubit systems with EPs.
@@ III.1 Non-interacting qubits
We begin our analysis of the spectral properties of the NHH in Eq. (1) with the simplest case, when $\textbf{J}\equiv\boldsymbol{0}$, and $f_{k}=1$, for $\forall k$. In this case the spectrum of $N$ non-interacting qubits is analogous to that of an $N$-dimensional hypercube [57, 58], namely
@@
which is generated by the spectrum of a single qubit $E^{\rm q}_{1,2}=\pm\alpha$. This single qubit energy spectrum is plotted in Fig. 1. In what follows, we assume that the energy levels are always arranged in ascending order, i.e.,
$E_{1}\leq E_{2}\leq\dots\leq E_{2N}.$ (7)
Figure 1: Energy spectrum of a single qubit, described by the NHH in Eq. (1), as a function of the applied magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$, according to Eqs. (6) and (3). The $x,y$ dependence of the energies reflects the $x,y$ dependence of the parameter $\alpha$. The purely real-valued spectrum consists of two intertwined Riemann sheets with an EP located at $(x=0,y=1)$. Moreover, at this EP the spectrum of the NHH diverges, according to Eq. (9) (see the main text for details).
@@
where $\hat{R}=\exp\left(-i\pi\sigma_{x}/4\right)$ is a rotation operator. Because the pseudo-Hermiticity of the parameterized NHH does not allow spontaneous symmetry breaking, the left eigenvectors are simply the complex conjugates of the right eigenvectors [44].
We now proceed with a more detailed examination of the spectral properties of a single qubit, since its properties are directly reflected in Eq. (6). As can be seen from Fig. 1, the qubit spectrum is characterized by the presence of a singular point at $(x=0,y=1)$. Indeed, according to Eq. (3) we have
@@
where the sign at infinity depends on the direction of the limit approach to the singularity (see Fig. 1). Furthermore, one finds the limit for the inner product of the normalized eigenvectors at this point is
@@
Equation (10) implies that the qubit eigenvectors coalesce at this singular point. In other words, the singularity behaves like an EP, although it is not an EP in the usual sense that is encountered in non-Hermitian systems with finite spectra. This seeming difficulty can be handled using mathematical tools borrowed from the projective space framework [59]. Specifically, this singularity can be cast into an ordinary EP in some extended NHH, where the eigenvalues converge at the EP (see Appendix B for details).
@@
Nevertheless, for our purposes here, we can safely ignore the infinite nature of the spectrum at such EPs, because we are solely interested in the loop trajectories around the EPs, along which both the energy spectrum and the parameters controlling the system ($x,y,{\textbf{J}}$), or equivalently, the complex magnetic fields and the qubit-qubit interactions remain finite. These EPs may therefore be termed as ’twisted EPs,’ due to the nontrivial wrapping of the spectrum at these points that leads to the divergent behavior of the NHH eigenvalues. For simplicity, however, in the remaining technical sections of this paper we shall continue to refer to such singularities as EPs.
@@
When dynamically encircling the EP in Fig. 1 in the parameter space $(x,y)$, the eigenstates of a qubit can be interchanged adiabatically without invoking any non-adiabatic transitions [44]. This stems from the fact that the eigenvalues are everywhere real. It is important to appreciate however, that while a winding trajectory around the EP in the space $(x,y)$ can have the form of simple loops in a plane, the actual orbits in the vector space (${\mathbb{C}}^{2}$ or ${\mathbb{R}}^{4}$) of the magnetic field $\vec{\mathbb{B}}_{||}$ are the loops on a curved hyperboloid.
@@
We note that the spectrum described above differs drastically from that studied in Ref. [44], where the qubit eigenvalues have a well-defined limit in Eq. (9). The key distinction lies in the fact that, unlike in Ref. [44], the Riemann energy sheets are now well separated at the special point $(x=0,y=0)$ where the NHH becomes Hermitian (see Fig. 1). This is a very useful property, since it can allow one to perform state transfer protocols over multiqubit states which are often generated and exploited in closed systems. This key property of the NHH of Eq. (1) for non-interacting qubits thus enables the combination and extension of existing quantum information protocols, originally devised for Hermitian systems, to non-Hermitian qubit systems as well. In the next subsection we extend this to the case of interacting qubits.
@@
Figure 2: Eigenspectrum of the NHH $\hat{H}$ in Eq. (1) as a function of the magnetic field $\vec{\mathbb{B}}_{||}^{i}(x,y)$ for (a) two interacting qubits and (b) three interacting qubits. The system parameters are (a) $f_{2}=2f_{1}=2$, $J_{12}=1$; and (b) $f_{1,2,3}=1$, $J_{12}=0.5$, $J_{13}=1$; $J_{23}=1.5$. The spectrum consists of a stack of Riemann sheet pairs, whose number grows with the number of interacting qubits in the system. The set of EPs, constituting the exceptional line (EL), is located at the same point $(x=0,y=1)$ as in the case of a single non-interacting qubit in Fig. 1. We note that the EPs are characterized by the same divergence behavior as in Fig. 1, in accordance with Eq. (9). As noted in the text, this does not affect the dynamics of trajectories around the EPs so, for the sake of clarity, we omit explicit depiction of the divergence here and in the figures that follow.